# Benchmarking Large Vision-Language Models on Fine-Grained Image Tasks: A Comprehensive Evaluation

### v2

@@ Abstract
Recent advancements in Large Vision-Language Models (LVLMs) have demonstrated remarkable multimodal perception capabilities, garnering significant attention. While numerous evaluation studies have emerged, assessing LVLMs both holistically and on specialized tasks, fine-grained image tasks—fundamental to computer vision—remain largely unexplored. To fill this gap, we introduce a comprehensive fine-grained evaluation benchmark, i.e.FG-BMK, comprising 1.01 million questions and 0.33 million images. Our evaluation systematically examines LVLMs from both human-oriented and machine-oriented perspectives, focusing on their semantic recognition and fine-grained feature representation capabilities. Through extensive experiments on twelve representative LVLMs/VLMs, we uncover key findings regarding the influence of training paradigms, modality alignment, perturbation susceptibility, and fine-grained category reasoning on task performance. This work provides critical insights into the limitations of current LVLMs and offers guidance for future data construction and model design in the development of more advanced LVLMs. Our code is open-source and available at https://anonymous.4open.science/r/FG-BMK-7B51.
@@ 1 Introduction
Large language models have recently achieved remarkable advancements, with models like GPT-4 [1] surpassing human-level performance across diverse tasks. Building on this, Large Vision-Language Models (LVLMs) have rapidly evolved, with models such as GPT-4o [1], Qwen [2], InternVL [8], and LLaVA-1.5 [24] demonstrating strong multimodal reasoning and perception capabilities.
In response to these advancements, various holistic and specialized evaluations have emerged to assess the capabilities of LVLMs. For instance, LVLM-eHub [48] and MMBench [51] provide broad evaluations of multimodal perception and reasoning, while specialized evaluations like DocVQA [31] and GQA [15] focus on specific tasks, such as document visual perception and visual reasoning. While these evaluations have provided valuable insights, a critical gap remains: there is no systematic evaluation of LVLMs on fine-grained image tasks, which focus on analyzing visual objects at the subordinate category level and are fundamental to computer vision [44]. Consequently, the capability boundaries of LVLMs in fine-grained tasks remain poorly understood.
To address this gap, we propose a comprehensive fine-grained evaluation benchmark, termed FG-BMK. This benchmark includes 1.01 million questions and 0.33 million images, providing a robust test bed for evaluating LVLMs. FG-BMK incorporates two evaluation paradigms: human-oriented and machine-oriented. The human-oriented evaluation employs dialogue-like interactions to assess a model’s ability to understand and respond to fine-grained visual queries in a conversational context. The machine-oriented evaluation focuses on two core fine-grained vision tasks—image retrieval and recognition—to directly measure the feature representation capabilities of LVLMs. Together, these evaluations enable a comprehensive assessment of LVLMs’ fine-grained feature representation and semantic recognition abilities.
@@ 2 Related Work
Large Language Models (LLMs) like GPT-4 [1] have achieved significant advancements in text comprehension, reasoning, and generation. Building on this foundation, Large Vision-Language Models (LVLMs) have emerged, demonstrating remarkable reasoning and perception capabilities across various tasks. Efforts to enhance LVLM performance have taken diverse approaches. BLIP [22] leverages noisy web data with bootstrapped captions, while LLaVA [24] utilizes GPT-4-generated instruction-following data to improve visual-linguistic understanding. BLIP-2 [21] focuses on integrating frozen image and text encoders, and Qwen2.5-VL [2] employs dynamic-resolution strategy to preserve native image resolution. InternVL3 [55] adopts a unified pre-training approach over multimodal data to improve effectiveness. Despite their success across various tasks, these LVLMs’ capabilities in fine-grained visual tasks remain underexplored. Our work seeks to fill this gap by evaluating their performance in this challenging domain, providing deeper insights into their potential and limitations.
@@
While LVLMs have shown impressive performance, various benchmarks have been developed to evaluate their capabilities in different domains. Specialized benchmarks like ChartQA [30] for chart understanding, DocVQA [31] for document visual perception, and GQA [15] for visual reasoning provide focused assessments. Additionally, evaluations in optical character recognition [25] and adversarial robustness [28] offer insights into specialized aspects of LVLM performance. Holistic evaluations, such as LVLM-eHub [48] and MMBench [51], assess general multimodal perception and reasoning, while others like MathVista [27] and MMMU [52] include expert-level problems spanning multiple disciplines.
However, comprehensive evaluations of LVLMs in fine-grained visual tasks remain scarce. Existing efforts, such as [12], primarily focus on fine-grained classification, and [54] evaluates classification and reasoning with limited questions. To address it, we propose FG-BMK, a benchmark designed to comprehensively evaluate LVLMs’ fine-grained feature representation and semantic recognition capabilities across diverse visual tasks.
@@
Fine-grained visual tasks [44, 49, 45, 47, 16, 53] are pivotal in applications like biodiversity monitoring [17], object retrieval [35], and product recommendation [43]. While LVLMs such as GPT-4, InternVL, and Qwen excel in tasks like OCR, visual question answering, and image captioning, fine-grained tasks remain a significant challenge. These tasks demand models to identify subtle, discriminative patterns in images [37] and leverage expert knowledge from LLMs for precise responses. To address this, we introduce a comprehensive benchmark and conduct extensive experiments to evaluate LVLMs’ performance on fine-grained tasks. Our study systematically highlights their limitations and offers actionable insights for advancing model design and training.
@@ 3.1 Evaluation Tasks and Metrics
While, the machine-oriented evaluation encompasses two fundamental vision tasks [44]: image retrieval and image classification. To evaluate the feature representation ability of LVLMs, we measure the accuracy of their visual features on these tasks. Following the DINOv2 approach [33], we use mean Average Precision (mAP) for retrieval and Top-1 accuracy for classification.
@@ 3.2 Data Curation
To ensure the quality and diversity of the data, we source images for the FG-BMK benchmark from 12 well-established fine-grained datasets, which helps avoid issues of inconsistent image quality and the risk of mislabeling often found in web-sourced images [34].
@@ 4.1 Models under Evaluation
Given the diverse landscape of existing LVLMs, we select nine widely-used open-source LVLMs, two closed-source models (GPT-4o-1120 [1] and Gemini-2.0-flash [13]) and one purely visual model with varying training strategies, as shown in Table 1. To enable clear comparisons and isolate the factors influencing performance, we use earlier versions of models from each family in machine-oriented evaluation, where their distinct characteristics are more evident and less affected by additional tricks or complex modifications. Further details about the evaluated models can be found in Appendix B.
@@ 4.2 Human-oriented Evaluation
Table 1: Training strategies of the open-source evaluated models. “DINOv2" is a purely visual model. “Con" denotes contrastive loss, “Gen" generative loss, “Mat" image-text matching loss, “Rec" reconstruction loss used in BEiT3, and “Dis" distillation loss used in DINOv2.
Model Vision Encoder Loss Function Training Data
Con Gen Mat Rec Dis $<$ 100M 100M $\sim$ 1B $>$ 1B
InternVL3-7B [55] ViT-L
InternVL-Chat-V1.1 [8] ViT-6B
LLaVA-1.5-7B [24] ViT-L
Qwen2.5-VL-7B [3] ViT-600M
Qwen-VL-Chat [2] ViT-G
BLIP-2-FLAN-T5-XL [21] ViT-G
EVA-CLIP [38] ViT-L
BEiT3 [42] ViT-L
CoCa [50] ViT-L
DINOv2 [33] ViT-L
@@
Figure 2: Results of InternVL3 [55] on true/false and multiple-choice questions across different levels of granularity on the CUB-200-2011 [41] dataset. The $x$-axis denotes the granularity of the recognition questions.
As shown in Figure 2, taking InternVL3 [55] as an example, its accuracy in answering true/false and multiple-choice questions declines as the granularity becomes finer. At the class level (e.g., “Is the class of the object in this image an Insecta/Aves?”), the model achieves 99.76% accuracy on multiple-choice questions and 99.77% on true/false questions. [^1] When questions are relatively simple, LVLMs achieve very high accuracy. The higher accuracy on multiple-choice questions compared to true/false questions may result from randomness. However, as the granularity narrows to the genus level, where negative categories are drawn from different genera within the same class (e.g., “Is the object in this image an albatross or a gull?”), the model’s accuracy on multiple-choice questions drops to 90.75%, a decrease of 9.01%. When the granularity is further refined to the species level, where negative categories are from different species within the same genus (e.g., “Is the object in this image a black-footed albatross/Laysan albatross?”), the accuracy drops further to 62.48% on true/false questions and 61.18% on multiple-choice questions. This demonstrates that the model struggles to distinguish between closely related species. The results for other LVLMs exhibit similar trends to those observed for InternVL3. Additional examples of multiple-choice and true/false questions can be found in Appendix C.1.
@@
To examine whether LVLMs exhibit knowledge bias in recognizing different fine-grained categories, we ranked their accuracy in answering true/false questions for each fine-grained category. As shown in Figure 3, using LLaVA [24] as an example, the model demonstrates highly inconsistent recognition abilities across categories, achieving nearly 90% accuracy for some while dropping to approximately 30% for others.
@@
Figure 3: Comparison of the original and fine-tuned LLaVA models on occurrence-balanced fine-grained bird categories. True/false question accuracy for each category is ranked, with blue dots representing the original model and yellow dots the fine-tuned model.
@@
Table 2: Attribute recognition accuracy of InternVL3 [55] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
@@
As shown in Table 2 of the paper and Table 12 in the appendix, LVLMs exhibit relatively stronger recognition abilities for attributes like pattern, size and length compared to others. For example, InternVL3 [55] and Qwen2.5-VL [3] achieve 50.13% and 45.12% average accuracy for pattern recognition, but only 30.95% and 29.30% for shape recognition. Notably, only a few attributes achieved an accuracy above 70%, and several attributes scored as low as 10%, highlighting significant room for improvement in LVLMs’ fine-grained attribute recognition.
We also observe model-specific differences in attribute recognition. For instance, while InternVL3 struggles more with pattern recognition compared to size, Gemini-2.0-flash [13] shows the opposite trend. Additionally, we find that the latest LVLMs show significant improvements in pattern and length recognition, while their advancements in color and shape recognition are more limited. Detailed results of the attribute recognition task are provided in Appendix C.3.
@@
Table 3: Comparison of LVLMs and fine-grained tailored models on classification tasks. “SA" denotes LVLM results fine-tuned on fine-grained datasets for short-answer questions, “LC" represents linear classifier results using LVLM visual features, and “FG-Tailored" refers to state-of-the-art fine-grained tailored model results.
@@
CUB-200-2011 [41] 85.60 91.65 93.10 [11]
Stanford Dog [18] 86.49 90.50 97.30 [6]
Stanford Car [19] 90.55 94.30 97.10 [23]
Food-101 [7] 95.25 95.67 98.60 [5]
FGVC Aircraft [29] 66.19 78.88 95.40 [36]
In Table 3, we compare the recognition accuracy of LVLMs with that of fine-grained tailored models. While LVLMs demonstrate notable performance in fine-grained recognition, their accuracy—whether in short-answer questions or using the linear classifier method—falls short of that achieved by fine-grained tailored models. This disparity may stem from the fact that LVLMs are primarily optimized for general tasks, with less emphasis placed on fine-grained domain capabilities. Enhancing their performance in specialized areas (e.g., fine-grained recognition) while preserving their strengths in general tasks presents an important and promising direction for future research.
@@ 4.3 Machine-oriented Evaluation
Compared to human-oriented evaluation, machine-oriented evaluation takes a more direct approach to assess the visual feature representations of LVLMs by employing two fundamental visual tasks: image retrieval and image recognition. This evaluation focuses on two key aspects: 1) Discriminability, which ensures the features can effectively distinguish between fine-grained categories, and 2) Robustness, which examines whether classification accuracy is maintained when visual features are subjected to perturbations.
To evaluate discriminability, we start by analyzing the performance of visual features on retrieval and classification tasks, following the setting of DINOv2 [33]. Figures 5 and 5 display retrieval and classification results across twelve fine-grained datasets, while Figures 7 and 7 provide the Friedman test results for these tasks.
Then, we increase the classification difficulty by merging fine-grained categories from different super-categories into a unified training and testing set. Figure 8 shows the classification accuracy both within a single super-category and across multiple meta-categories. Visual features with strong discriminability should demonstrate high accuracy even when faced with diverse data sources. Furthermore, we examine how the size of the vision encoder in LVLMs affects classification accuracy, as illustrated in Figure 9.
Alignment, a critical component of LVLMs, bridges the gap between vision and language modalities. To explore its influence on visual features, we compare the classification accuracy of original visual features with those aligned to textual features in fine-grained classification tasks, as shown in Table 4.
To investigate robustness, we introduce perturbations to the visual features using projected gradient descent [28] and analyze their impact on classification accuracy. The results of this experiment are presented in Table 5.
@@
In Figure 5 and Figure 5, vision encoders of LVLMs trained with contrastive paradigms (e.g., EVA-CLIP, InternVL, DINOv2) outperform those trained with reconstruction (BEiT3) and generative paradigms (Qwen) on fine-grained retrieval and classification tasks. As shown in Figure 7 and Figure 7, the Nemenyi statistical test results also indicate that InternVL, EVA-CLIP and DINOv2 perform significantly better than Qwen and BEiT3. In the multiple meta-category classification results (cf. Figure 8), EVA-CLIP maintains high accuracy, with an average decrease of just 1.96% compared to the single meta-category setting, while Qwen and BEiT3 show larger drops of 4.16% and 7.41%, respectively. This underscores the effectiveness of contrastive training paradigms in fine-grained tasks.
@@
Figure 8: Classification results of LVLM visual features on fine-grained datasets. “Single" denotes accuracy from training on a single meta-category, while “Multiple" reflects accuracy from training on a unified dataset combining multiple meta-categories.
@@
Table 4: Classification results of LLaVA visual features on fine-grained datasets before and after alignment with textual features. “Origin" refers to results from the original visual features extracted by the vision encoder. “Aligned" indicates results after alignment with text features using data with inconsistent granularity. “Aligned-FG" shows results after alignment with data of matching visual-text granularity.
@@
Alignment plays a critical role in LVLMs, bridging the gap between visual and textual features. To investigate its impact on fine-grained visual feature discriminability, we compare the classification accuracy of LLaVA’s [24] original visual features with those aligned to textual features on fine-grained classification tasks. As shown in the first two columns of Table 4, the original visual features demonstrate superior classification capabilities, outperforming the aligned features by an average of 3.39%.
@@
To validate the impact of granularity inconsistency, we construct a new alignment dataset where the textual descriptions match the granularity of the fine-grained objects in images. We then retrain the alignment module in LLaVA using this dataset. As shown in Table 4, the classification accuracy of visual features aligned with fine-grained textual content improves significantly, with an increase of 2.55% on Stanford Dogs and 1.73% on Stanford Cars. While the aligned visual features still underperform compared to the original features, these results underscore the detrimental effects of granularity inconsistency on visual feature discriminability and highlight the importance of ensuring granularity alignment between image and textual content.
@@
Table 5: Classification results of LVLMs’ original and perturbed visual features on the fine-grained dataset CUB-200-2011 and the generic dataset CIFAR-100. “Origin" refers to results with original features, while “Perturbed" indicates results with perturbed features.
@@
As shown in Table 5, perturbing the visual features of EVA-CLIP causes a significant drop in classification accuracy on the fine-grained dataset CUB-200-2011 (from 88.95% to 24.94%). In contrast, on the generic dataset CIFAR-100 [20], the accuracy declines more modestly (from 93.05% to 50.76%). Similar trends are observed for CoCa and DINOv2, indicating that fine-grained tasks are more susceptible to perturbations than generic tasks. We attribute this vulnerability to the coarse-grained and noisy data used during training, which limits the discriminability of features across fine-grained categories, making it challenging to distinguish between them when visual features are perturbed.
In contrast, the Vision Transformer [46] trained on the ImageNet [10] dataset with cross-entropy loss demonstrates superior robustness to perturbations. It shows only minor declines in classification accuracy on both fine-grained and generic datasets. This suggests that adopting alternative training paradigms or incorporating high-quality, fine-grained data (as seen in ImageNet) during training could help improve the robustness of visual features in LVLMs.

### v3

@@ Abstract
Recent advancements in Large Vision-Language Models (LVLMs) have demonstrated remarkable multimodal perception capabilities, garnering significant attention. While numerous evaluation studies have emerged, assessing LVLMs both holistically and on specialized tasks, fine-grained image tasks—fundamental to computer vision—remain largely unexplored. To fill this gap, we introduce a comprehensive fine-grained evaluation benchmark, i.e., FG-BMK, comprising 1.01 million questions and 0.28 million images. Our evaluation systematically examines LVLMs from both human-oriented and machine-oriented perspectives, focusing on their semantic recognition and fine-grained feature representation capabilities. Through extensive experiments on twelve representative LVLMs/VLMs, we uncover key findings regarding the influence of training paradigms, modality alignment, perturbation susceptibility, and fine-grained category reasoning on task performance. This work provides critical insights into the limitations of current LVLMs and offers guidance for future data construction and model design in the development of more advanced LVLMs. Our code is open-source and available at https://github.com/SEU-VIPGroup/FG-BMK.
@@ 1 Introduction
Large language models have recently achieved remarkable advancements, with models like GPT-4 (OpenAI, 2023) surpassing human-level performance across diverse tasks. Building on this, Large Vision-Language Models (LVLMs) have rapidly evolved, with models such as GPT-4o (OpenAI, 2023), Qwen (Bai et al., 2023), InternVL (Chen et al., 2024), and LLaVA-1.5 (Liu et al., 2024a) demonstrating strong multimodal reasoning and perception capabilities.
In response to these advancements, various holistic and specialized evaluations have emerged to assess the capabilities of LVLMs. For instance, LVLM-eHub (Xu et al., 2025) and MMBench (Yuan et al., 2024) provide broad evaluations of multimodal perception and reasoning, while specialized evaluations like DocVQA (Mathew et al., 2021) and GQA (Hudson and Manning, 2019) focus on specific tasks, such as document visual perception and visual reasoning. While these evaluations have provided valuable insights, a critical gap remains: there is no systematic evaluation of LVLMs on fine-grained image tasks, which focus on analyzing visual objects at the subordinate category level and are fundamental to computer vision (Wei et al., 2022b). Consequently, the capability boundaries of LVLMs in fine-grained tasks remain poorly understood.
To address this gap, we propose a comprehensive fine-grained evaluation benchmark, termed FG-BMK. This benchmark includes 1.01 million questions and 0.28 million images, providing a robust test bed for evaluating LVLMs. FG-BMK incorporates two evaluation paradigms: human-oriented and machine-oriented. The human-oriented evaluation employs dialogue-like interactions to assess a model’s ability to understand and respond to fine-grained visual queries in a conversational context. The machine-oriented evaluation focuses on two core fine-grained vision tasks—image retrieval and recognition—to directly measure the feature representation capabilities of LVLMs. Together, these evaluations enable a comprehensive assessment of LVLMs’ fine-grained feature representation and semantic recognition abilities.
@@ 2 Related Work
Large Language Models (LLMs) like GPT-4 (OpenAI, 2023) have made notable strides in text comprehension, reasoning, and generation. Building on this foundation, Large Vision-Language Models (LVLMs) have emerged with impressive reasoning and perception abilities across diverse tasks. Enhancing LVLM performance has taken various directions: BLIP (Li et al., 2022) uses noisy web data with bootstrapped captions; LLaVA (Liu et al., 2024a) employs GPT-4-generated instruction-following data; BLIP-2 (Li et al., 2023) integrates frozen image and text encoders; Qwen2.5-VL (Bai et al., 2023) leverages dynamic-resolution strategy for preserving native image resolution; InternVL3 (Zhu et al., 2025) unifies pre-training over multimodal data for improved effectiveness. While achieving success on many tasks, fine-grained visual challenges for these LVLMs remain underexplored. This work addresses this gap by assessing their performance in fine-grained domains, uncovering both their potential and limitations.
@@
While LVLMs have shown impressive performance, various benchmarks have been developed to evaluate their capabilities in different domains. Specialized benchmarks like ChartQA (Masry et al., 2022) for chart understanding, GQA (Hudson and Manning, 2019) for visual reasoning, and CAPability (Liu et al., 2025) for caption evaluation provide focused assessments. Additionally, evaluations in optical character recognition (Liu et al., 2024c) and adversarial robustness (Madry et al., 2018) offer insights into specialized aspects of LVLM performance. Holistic evaluations, such as LVLM-eHub (Xu et al., 2025) and MMBench (Yuan et al., 2024), assess general multimodal perception and reasoning, while others like MathVista (Lu et al., 2024) and MMMU (Yue et al., 2024) include expert-level problems spanning multiple disciplines.
However, comprehensive evaluations of LVLMs in fine-grained visual tasks remain scarce. Existing efforts, such as (Geigle et al., 2024), primarily focus on fine-grained classification, and (Zhang et al., 2024c) evaluates classification and reasoning with limited questions. To address it, we propose FG-BMK, a benchmark designed to comprehensively evaluate LVLMs’ fine-grained feature representation and semantic recognition capabilities across diverse visual tasks.
@@
Fine-grained visual tasks (Wei et al., 2022b; Xu et al., 2022; Wei et al., 2024; Zhang et al., 2024a; Jing et al., 2024a; Wang et al., 2024; Liu et al., 2024b; Zhang et al., 2024b) are pivotal in applications like biodiversity monitoring (Jing et al., 2024b), object retrieval (Shen et al., 2022), and product recommendation (Wei et al., 2022a). While LVLMs such as GPT-4, InternVL, and Qwen excel in tasks like OCR, visual question answering, and image captioning, fine-grained tasks remain a significant challenge. These tasks demand models to identify subtle, discriminative patterns in images (Song et al., 2020) and leverage expert knowledge from LLMs for precise responses. To address this, we introduce a comprehensive benchmark and conduct extensive experiments to evaluate LVLMs’ performance on fine-grained tasks. Our study systematically highlights their limitations and offers actionable insights for advancing model design and training.
@@ 3.1 Evaluation Tasks and Metrics
While, the machine-oriented evaluation encompasses two fundamental vision tasks (Wei et al., 2022b): image retrieval and image classification. To evaluate the feature representation ability of LVLMs, we measure the accuracy of their visual features on these tasks. Following the DINOv2 approach (Oquab et al., 2023), we use mean Average Precision (mAP) for retrieval and Top-1 accuracy for classification.
@@ 3.2 Data Curation
To ensure the quality and diversity of the data, we source images for the FG-BMK benchmark from 12 well-established fine-grained datasets, which helps avoid issues of inconsistent image quality and the risk of mislabeling often found in web-sourced images (Radford et al., 2021).
@@ 4.1 Models under Evaluation
Given the diverse landscape of existing LVLMs, we select nine widely-used open-source LVLMs, two closed-source models (GPT-4o-1120 (OpenAI, 2023) and Gemini-2.0-flash (Gemini Team, 2024)) and one purely visual model with varying training strategies, as shown in Table 1. To enable clear comparisons and isolate the factors influencing performance, we use earlier versions of models from each family in machine-oriented evaluation, where their distinct characteristics are more evident and less affected by additional tricks or complex modifications. Further details about the evaluated models can be found in Appendix B.
@@ 4.2 Human-oriented Evaluation
Table 1: Training strategies of the open-source evaluated models. “DINOv2” is a purely visual model. “Con” denotes contrastive loss, “Gen” generative loss, “Mat” image-text matching loss, “Rec” reconstruction loss used in BEiT3, and “Dis” distillation loss used in DINOv2.
Model Vision Size Loss Function Training Data
Con Gen Mat Rec Dis << 0.1B 0.1B $\sim$ 1B >> 1B
InternVL3-7B (Zhu et al., 2025) ViT-L
InternVL-Chat (Chen et al., 2024) ViT-6B
LLaVA-1.5-7B (Liu et al., 2024a) ViT-L
Qwen2.5-VL-7B (Bai et al., 2025) ViT-600M
Qwen-VL-Chat (Bai et al., 2023) ViT-G
BLIP-2-XL (Li et al., 2023) ViT-G
EVA-CLIP (Sun et al., 2023) ViT-L
BEiT3 (Wang et al., 2023) ViT-L
CoCa (Yu et al., 2022) ViT-L
DINOv2 (Oquab et al., 2023) ViT-L
@@
Figure 2: Results of InternVL3 (Zhu et al., 2025) on true/false and multiple-choice questions across different levels of granularity on the CUB-200-2011 (Wah et al., 2011) dataset. The $x$-axis denotes the granularity of the recognition questions.
As shown in Figure 2, taking InternVL3 (Zhu et al., 2025) as an example, its accuracy in answering true/false and multiple-choice questions declines as the granularity becomes finer. At the class level (e.g., “Is the class of the object in this image an Insecta/Aves?”), the model achieves 99.76% accuracy on multiple-choice questions and 99.77% on true/false questions. [^1] When questions are relatively simple, LVLMs achieve very high accuracy. The higher accuracy on multiple-choice questions compared to true/false questions may result from randomness. However, as the granularity narrows to the genus level, where negative categories are drawn from different genera within the same class (e.g., “Is the object in this image an albatross or a gull?”), the model’s accuracy on multiple-choice questions drops to 90.75%, a decrease of 9.01%. When the granularity is further refined to the species level, where negative categories are from different species within the same genus (e.g., “Is the object in this image a black-footed albatross/Laysan albatross?”), the accuracy drops further to 62.48% on true/false questions and 61.18% on multiple-choice questions. This demonstrates that the model struggles to distinguish between closely related species. The results for other LVLMs exhibit similar trends to those for InternVL3. Additional examples of multiple-choice and true/false questions can be found in Appendix C.1.
@@
To examine whether LVLMs exhibit knowledge bias in recognizing different fine-grained categories, we ranked their accuracy in answering true/false questions for each fine-grained category. As shown in Figure 3, using LLaVA (Liu et al., 2024a) as an example, the model demonstrates highly inconsistent recognition abilities across categories, achieving nearly 90% accuracy for some while dropping to approximately 30% for others.
Figure 3: Comparison of the original (blue dots) and fine-tuned (yellow dots) LLaVA models on occurrence-balanced fine-grained bird categories. True/false accuracy per category is ranked.
@@
Table 2: Attribute recognition accuracy of InternVL3 (Zhu et al., 2025) on the CUB-200-2011 (Wah et al., 2011) dataset (values in parentheses represent the average accuracy for each attribute).
@@
As shown in Table 2 of the paper and Table 14 in the appendix, LVLMs exhibit relatively stronger recognition abilities for attributes like pattern, size and length compared to others. For example, InternVL3 (Zhu et al., 2025) and Qwen2.5-VL (Bai et al., 2025) achieve 50.13% and 45.12% average accuracy for pattern recognition, but only 30.95% and 29.30% for shape recognition. Notably, only a few attributes achieved an accuracy above 70%, and several attributes scored as low as 10%, highlighting significant room for improvement in LVLMs’ fine-grained attribute recognition.
We also observe model-specific differences in attribute recognition. For instance, while InternVL3 struggles more with pattern recognition compared to size, Gemini-2.0-flash (Gemini Team, 2024) shows the opposite trend. Additionally, we find that the latest LVLMs show significant improvements in pattern and length recognition, while their advancements in color and shape recognition are more limited. Detailed results of the attribute recognition task are provided in Appendix C.3.
@@
In Table 3, we compare the recognition accuracy of LVLMs with that of fine-grained tailored models. While LVLMs demonstrate notable performance in fine-grained recognition, their accuracy—whether in short-answer questions or using the linear classifier method—falls short of that achieved by fine-grained tailored models. This disparity may stem from the fact that LVLMs are primarily optimized for general tasks, with less emphasis placed on fine-grained domain capabilities (e.g., CAP (Behera et al., 2021) uses context-aware attentional pooling to capture hierarchical contextual information from pixel to region to image level, leading to a 1.53% improvement in classification accuracy).
Table 3: Comparison of LVLMs and fine-grained tailored models on classification tasks. “SA” denotes LVLM results fine-tuned on fine-grained datasets for short-answer questions, “LC” represents linear classifier results using LVLM visual features, and “FG-Tailored” refers to state-of-the-art fine-grained tailored model results.
@@
CUB-200-2011 85.60 91.65 93.10 (Diao et al., 2022)
Stanford Dog 86.49 90.50 97.30 (Bera et al., 2022)
Stanford Car 90.55 94.30 97.10 (Liu, 2024)
Food-101 95.25 95.67 98.60 (Behera et al., 2021)
FGVC Aircraft 66.19 78.88 95.40 (Sikdar et al., 2024)
Due to the architectural paradigm of LVLMs (typically ViT + MLP + LLM), these components cannot be directly applied. However, the core idea of attending to hierarchical-level details can still be incorporated into LVLMs (e.g., InternVL3 (Zhu et al., 2025) not only takes the full image as input, but also divides the image into parts to obtain more fine-grained visual representations). Enhancing LVLMs performance in specialized areas (e.g., fine-grained recognition) while preserving their strengths in general tasks presents an important and promising direction for future research.
@@ 4.3 Machine-oriented Evaluation
Compared to human-oriented evaluation, machine-oriented evaluation takes a more direct approach to assess the visual feature representations of LVLMs by employing two fundamental visual tasks: image retrieval and image recognition. This evaluation focuses on two key aspects: 1) Discriminability—the ability of visual features to distinguish fine-grained categories, and 2) Robustness—maintaining accuracy under feature perturbations.
To evaluate discriminability, we start by analyzing the performance of visual features on retrieval and classification tasks, following the setting of DINOv2 (Oquab et al., 2023). Figures 5 and 5 display retrieval and classification results across twelve fine-grained datasets, while Figures 7 and 7 provide the Friedman test results for these tasks.
@@
We then increase classification difficulty by merging fine-grained categories from different super-categories into a unified training and testing set. Figure 8 presents classification accuracy within a single super-category and across multiple meta-categories. Visual features with strong discriminability should maintain high accuracy even with diverse data sources. Additionally, we analyze how the vision encoder size in LVLMs impacts classification accuracy, as shown in Figure 9.
Furthermore, we examine how vision-language alignment—a critical component of LVLMs—affects visual feature quality. Specifically, we compare classification accuracy between original visual features and those aligned to textual features in fine-grained classification tasks, as shown in Table 4.
To investigate robustness, we introduce perturbations to visual features using projected gradient descent (Madry et al., 2018) and analyze their impact on classification accuracy. Results of this experiment are presented in Table 5.
@@
In Figure 5 and Figure 5, vision encoders trained with contrastive paradigms (e.g., EVA-CLIP, InternVL, DINOv2) outperform those trained with reconstruction (BEiT3) and generative paradigms (Qwen) on fine-grained retrieval and classification tasks. As shown in Figure 7 and Figure 7, Nemenyi test results also indicate that InternVL, EVA-CLIP, and DINOv2 perform significantly better than Qwen and BEiT3. In multi meta-category classification (cf. Figure 8), EVA-CLIP maintains high accuracy, with an average drop of just 1.96% compared to the single-category setting, while Qwen and BEiT3 show larger drops of 4.16% and 7.41%, respectively. This underscores the effectiveness of contrastive paradigms in fine-grained tasks.
@@
Figure 8: Classification results of LVLM visual features on fine-grained datasets. “Single” denotes accuracy from training on a single meta-category, while “Multiple” reflects accuracy from training on a unified dataset combining multiple meta-categories.
@@
Table 4: Classification accuracy of LLaVA visual features before and after alignment. “Origin” denotes original features from the vision encoder. “Aligned” denotes features aligned to text with inconsistent granularity, while “Aligned-FG” denotes those aligned to fine-grained text.
@@
Alignment plays a critical role in LVLMs by bridging the gap between visual and textual features. To assess its impact on fine-grained visual feature discriminability, we compare the classification accuracy of LLaVA’s (Liu et al., 2024a) original visual features with those aligned to textual features on fine-grained classification tasks. As shown in the first two columns of Table 4, the original features demonstrate superior classification performance, outperforming the aligned ones by an average of 3.39%.
@@
To validate the impact of granularity inconsistency, we construct a new alignment dataset where the textual descriptions match the granularity of the fine-grained objects in images. We then retrain the alignment module in LLaVA using this dataset. As shown in Table 4, the classification accuracy of visual features aligned with fine-grained textual content improves significantly, with an increase of 2.55% on Stanford Dogs and 1.73% on Stanford Cars. While the aligned visual features still underperform compared to the original features, these results underscore the detrimental effects of granularity inconsistency on visual feature discriminability and highlight the importance of ensuring granularity alignment between image and textual content. Further experiments in Appendix D.2 reveal that enhancing the fine-grained discriminability of visual features during the alignment stage can lead to improved LVLM performance on fine-grained tasks.
@@
Table 5: Classification results of LVLMs’ original and perturbed visual features on the fine-grained dataset CUB-200-2011 and the generic dataset CIFAR-100. “Origin” refers to results with original features, while “Perturbed” indicates results with perturbed features.
@@
As shown in Table 5, perturbing the visual features of EVA-CLIP causes a significant drop in classification accuracy on the fine-grained dataset CUB-200-2011 (from 88.95% to 24.94%). In contrast, on the generic dataset CIFAR-100 (Krizhevsky and Hinton, 2009), the accuracy declines more modestly (from 93.05% to 50.76%). Similar trends are observed for CoCa and DINOv2, indicating that fine-grained tasks are more susceptible to perturbations than generic tasks. We attribute this vulnerability to coarse-grained, noisy training data, which limits feature discriminability across fine-grained categories, making it hard to distinguish them when visual features are perturbed.
In contrast, the Vision Transformer (Wightman, 2019) trained on the ImageNet (Deng et al., 2009) dataset with cross-entropy loss demonstrates superior robustness to perturbations. It shows only minor declines in classification accuracy on both fine-grained and generic datasets. This suggests that adopting alternative training paradigms or incorporating high-quality, fine-grained data (as seen in ImageNet) during training could help improve the robustness of visual features in LVLMs.